# Distribution is Unlimited. PA# AFRL-2024-2340

### v1

@@ Abstract
This work explores the conditions under which global contraction manifests in the leaky continuous time reservoirs, thus guaranteeing Generalized Synchronization. Results on continuous time reservoirs make use of the logarithmic norm of the connectivity matrix. Further analysis yields some simple guidelines on how to better construct the connectivity matrix in these systems. Additionally, we outline how the Universal Approximation Property of discrete time reservoirs is readily satisfied by virtue of the activation function being contracting, and how continuous time reservoirs may inherit a limited form of universal approximation by virtue of them overlapping with Neural Ordinary Differential Equations. The ability of the Reservoir Computing framework to universally approximate topological conjugates, along with their fast training, make them a compelling data-driven, black-box surrogate of dynamical systems, and a potential mechanism for developing digital twins.
@@ I Introduction
Echo State Networks (ESN) and Liquid State Machines (LSM) were proposed as promising architectures for neural networks more than two decades ago. The ESN approach comes from the perspective of artificial neural networks using a discrete time formulation and sigmoid activation function [1]. The LSM approach, on the other hand, comes (more directly) from the perspective of biological neural networks, where the network vector field follows the integrate-and-fire neuron model in continuous time [2]. Soon after, the term “reservoir” was first invoked in Steil [3], and the term Reservoir Computing (RC) was coined in Verstraeten et al. [4] while attempting to implement these architectures in field-programmable gated arrays.
Before continuing further, we define some terms to avoid potential confusion. We use the term Reservoir Computing to mean the series of steps used to set up and train a Recurrent Neural Network (RNN), which is treated as a driven dynamical system. The term reservoir refers to the driven state-space model; it has fixed or prescribed structure and has not yet been trained. The trained or autonomous reservoir refers to the autonomous state-space model, which incorporates the trained structure from a readout “layer”. As such, RC is a framework that utilizes the dynamics of an arbitrarily prescribed reservoir, where the training is performed only on the readout “layer”. The dynamics of the reservoir may be linear or nonlinear, but the corresponding readout must then be nonlinear and linear respectively.
The RC framework results in RNNs with some extremely promising aspects. It is able to predict autonomous dynamical systems with higher accuracy and smaller network sizes than deep feedforward networks and Long-Short Term Memory (LSTM) recurrent networks [5]. The training process is non-iterative thus quick – if the dynamics are nonlinear, the training requires only a linear fit, compounding on the aforementioned benefits of RC. This is therefore a promising framework, with compelling reasons to understand its theoretical underpinnings and the source of its many strengths. This has yielded considerable research efforts since its conception, particularly on the role of contraction in the listening phase of RC [6, 7, 8, 9, 10].
@@
The recognition that GS and the ESP were closely related phenomena was a particularly fruitful realization [18, 19, 20, 21, 22] and one of the core motivations of this research effort. In this work, we attempt to connect yet another concept – contraction theory [23, 24, 25, 26] – with GS and the ESP. The ESP in continuous time reservoirs is relatively understudied compared to their discrete time counterparts. We give a preamble with discrete time reservoirs and ESP, before pivoting to continuous time reservoirs. As far as the authors are aware of, Hart [20] provides the only results on the sufficient conditions of continuous time ESP. Our approach differs from that of Hart [20] but yields similar results that can be quickly computed.
The contraction that we invoke is better described as globally uniform state-contraction. Even though locally uniform state-contraction is sufficient for the ESP, the claims that we set out to make are nontrivial when considering all possible local cases, so we resort to the narrower global case. From here on, the prefix globally uniform state- is always implied when mentioning contraction, unless otherwise specified. It should also be understood that all inputs $u$ to the reservoir are measurements of an autonomous dynamical system with invertible velocity field and bounded trajectories.
It is important to keep in mind that ESP and GS are only necessary conditions for the successful implementation of the RC framework; it guarantees very little about the resulting autonomous reservoir and the quality of its predictions. Even the Universal Approximation Property (UAP) is not necessary, but a welcomed property – when possible – for a successfully trained reservoir to be a universal predictor of autonomous dynamical systems.
The organization of this work is as follows. Section II goes over the basic structure of discrete time reservoirs. It gives the current understanding of the role of contraction and displays how certain contracting properties can be leveraged. Section III handles contraction in continuous time reservoirs, which bears significant resemblance to the approach in Section II. Our results provide a quickly computable sufficient condition that can be easily adjusted to guarantee contraction. Section IV discusses the UAP in the context of contracting activation functions, and topological conjugacy of the autonomous reservoir. Lastly, to avoid lengthy tangents, the appendices cover some specific details and claims made in the main sections.
@@ II Discrete Time Systems
$x_{t+1}=\sigma(Ax_{t}+Bu_{t})$ (1)
$y_{t+1}=\sigma(Ay_{t}+Bu_{t})$
These driven dynamical systems described in (1) are called reservoirs with $N$ nodes in the context of this work. The connectivity matrix $A\in\mathbb{R}^{N\times N}$ describes the connections between the reservoir nodes. Matrix $B\in\mathbb{R}^{N\times M}$ is the input gain of the reservoir, designating how much each node is coupled to the inputs. The activation function, typically the hyperbolic tangent function, is the same for every node and is applied component-wise to its vector argument, i.e. $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{N}$ such that $\sigma([v_{1},v_{2}]^{\top})=[\sigma(v_{1}),\sigma(v_{2})]^{\top}$.
It is not expected that these two systems will reach a static equilibrium, since they are both driven by a time varying input $u$. Instead, consider defining a virtual displacement system (VDS) as $x_{t}-y_{t}$ to understand the long-term behavior of the individual systems. Specifically, it is sufficient that if the VDS shrinks according to some norm for every $t\geq 0$, the two parent systems in (1) will asymptotically approach the same solution and be identically synchronized, which implies that the the reservoir is synchronized (in the sense of GS) to the dynamical systems generating the inputs [14, 16].
$\norm{x_{t+1}-y_{t+1}}\leq k\norm{x_{t}-y_{t}}$ (2)
For some vector $v$, let $\norm{v}$ denote any vector norm unless otherwise specified. The property described by (2) yields exponential stability with constant $10\leq k<$ 1, which also guarantees asymptotic stability. To demonstrate exponential stability, we make use of the contraction mapping property of the sigmoid family of functions for any two vectors $a\neq b$ in $\mathbb{R}^{N}$.
$\norm{\sigma(a)-\sigma(b)}\leq\norm{a-b}$ (3)
In one dimension, contractions guarantee that two distinct points on its curve cannot generate a slope greater than unity and is equivalent to limiting the supremum magnitude of the derivative of $\sigma$ to unity. In higher dimensions, a contracting $\sigma$ can only shorten the distance between two vectors, even if a shift is applied. For some matrix $M$, let $\norm{M}$ denote any induced matrix norm. The matrix and vector norms throughout this work will be implied through consistent usage of upper- and lower-case letters of the respective arguments. Using (3), we show that the VDS is globally exponentially stable when driven by any bounded $u$, so long as the induced matrix norm $\norm{A}<1$. See Appendix A for additional details.
$\norm{\sigma(Ax_{t}+Bu_{t})-\sigma(Ay_{t}+Bu_{t})}$ $=\norm{A(x_{t}-y_{t})}$ (4)
@@
Defining in this context $\norm{A}_{\ell}$ as some arbitrary norm, a broader sufficient condition for stability is $\min_{\ell}\norm{A}_{\ell}<1$. For ease of calculation, $\ell$ is usually chosen to correspond to the $1$-, $2$-, or $\infty$-norm. We have shown that when $\norm{A}<1$, the systems in (1) will uniformly and exponentially approach one another at the rate no slower than $\norm{A}$ regardless of initial conditions. If $A$ can be prescribed, then a reservoir can be constructed such that the VDS is always globally exponentially stable.
$\norm{x_{t}-y_{t}}\leq\norm{A}^{t}\norm{x_{0}-y_{0}}\Longrightarrow\lim_{t\to\infty}{x_{t}-y_{t}}=0$ (5)
The rate at which the distance shrinks depend on the choice of norm. Also notice that the RHS of (4) is independent of $u_{t}$ and $B$. All this implies that a particular reservoir, defined by the triplet $(A,B,\sigma)$ with $\norm{A}<1$, will map every point of the input trajectory $u=\{u_{t}\}$ to a unique point in the reservoir trajectory $x=\{x_{t}\}$ in the infinite time limit [18, 14]. Necessarily, the ESP holds, and a unique GS occurs between the drive and response system.
$\lim_{t\to\infty}\frac{1}{t}\log\frac{\norm{x_{t}-y_{t}}}{\norm{x_{0}-y_{0}}}=\Lambda\leq\log\norm{A}<0$ (6)
This condition on $A$ is closely related to the largest conditional Lyapunov exponent (CLE) of the system $\Lambda$, conditioned on the driving signal $u$. In fact, $\log\norm{A}$ is the negative-definite upper bound of $\Lambda$, which is itself the upper bound on all the other CLEs. This implies GS, so the reservoir implicitly has a globally attractive trajectory $x=\{x_{t}\}$ unique to any given $u=\{u_{t}\}$ as $t\to\infty$. In the infinite time limit, there exists a unique continuous function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$, corresponding to the triplet $(A,B,\sigma)$ with constraint $\norm{A}\leq 1$, that maps the driving signal to the response, i.e. $x_{t}=\Phi(u_{t})$ as $t\to\infty$. For any desired error $\varepsilon>0$ and any chosen norm, there exist some function $\Phi$ and some time $T>0$ such that $\norm{x_{T}-\Phi(u_{T})}<\varepsilon$ [17, 27].
@@ III.1 Similarity to Discrete Time
In continuous time, the following situation is considered instead. Suppose that we are given some bounded and time varying input signal $u(t)\in\mathbb{R}^{M}$ in the time window $t\in[0,T]$ that is driving two identical reservoirs in the vector field according to the equations below.
$\dot{x}=-Cx+\sigma(Ax+Bu)$ (7)
$\dot{y}=-Cy+\sigma(Ay+Bu)$
As before, these two reservoirs have states $x(t)\in\mathbb{R}^{N}$ and $y(t)\in\mathbb{R}^{N}$ that stem from different initial conditions $x(0)\neq y(0)$. The matrices $A\in\mathbb{R}^{N\times N}$ and $B\in\mathbb{R}^{N\times M}$ serve the same purposes as the discrete time case, but an additional leak matrix $C\in\mathbb{R}^{N\times N}$ is necessary for stability, as will be demonstrated. The term leak is in reference to the leak current in biological neurons. For mathematical generality (as compared to biological realism), we impose that $C$ be symmetric positive definite (SPD). The activation function $\sigma$ is from the usual sigmoid family, applied component-wise as before. Define the VDS by $z(t)\equiv x(t)-y(t)$, with dynamics given by the following equation.
$\dot{z}=-C(x-y)+\sigma(Ax+Bu)-\sigma(Ay+Bu)$ (8)
@@
$\langle\dot{z},z\rangle=\underbrace{-\langle Cz,z\rangle}_{\leq-s_{N}(C)z^{2}}+\underbrace{\langle\sigma(Ax+Bu)-\sigma(Ay+Bu),x-y\rangle}_{\leq s_{1}(A)z^{2}}$ (9)
@@
$\dot{V}(\dot{z},z)\leq-2kV(z)\Longrightarrow\norm{z(t)}_{2}\leq\norm{z(0)}_{2}e^{-kt}$ (10)
Given the evolution of this Lyapunov function, $k>0$ results in the VDS being globally exponentially stable, making the parent systems (7) identically synchronized to one another, and the reservoir generally synchronized to the input. Since the evolution $\dot{V}(\dot{z},z)$ is ultimately a function of $z$ and $u$, $V(z)$, it is a control-Lyapunov function rather than a simple Lyapunov function. This is an important distinction as it suggests that $u$ plays a pivotal role in determining the attracting manifold, but plays no role determining whether the manifold is attracting. It is perhaps also a partial explanation as to why reservoirs robustly adapt to many input signals. Similar to the discrete time case, $-k$ is the upper bound of the largest CLE of the system.
$\lim_{t\to\infty}\frac{1}{t}\log\frac{\norm{z(t)}_{2}}{\norm{z(0)}_{2}}=\Lambda\leq-k<0$ (11)
In the asymptotic limit, the attracting trajectory of $x(t)$ is unique for a given input $u(t)$, up to the quadruplet $(A,B,C,\sigma)$ and the aforementioned constraints on $A$ and $C$. It is common in practice to use the identity matrix $I$ instead of the generic SPD matrix $C$. Of course, the identity matrix is itself SPD, but this fixes $s_{N}(I)=1$ and restricts $s_{1}(A)=\norm{A}_{2}<1$, displaying contraction with same triplet $(A,B,\sigma)$ as the 2-norm discrete time case, but with leaky integrate-and-fire dynamics. In the infinite time limit, there exists a continuous function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$ that maps $u(t)$ to $x(t)$, i.e. $x(t)=\Phi(u(t))$ as $t\to\infty$. Alternatively put, for any tolerance $\varepsilon>0$ and any chosen norm, there exist some function $\Phi$ and time $T>0$ such that $\norm{x(T)-\Phi(u(T))}_{2}<\varepsilon$.

### v2

@@ Abstract
This paper explores the conditions under which global contraction manifests in the leaky continuous time reservoirs, thus guaranteeing generalized synchronization. Results on continuous time reservoirs make use of the logarithmic norm of the connectivity matrix. Further analysis yields some simple guidelines on how to better construct the connectivity matrix in these systems. Additionally, we outline how the universal approximation property of discrete time reservoirs is readily satisfied by virtue of the activation function being contracting, and how continuous time reservoirs may inherit a limited form of universal approximation due to their overlap with neural ordinary differential equations. The ability of the reservoir computing framework to universally approximate topological conjugates, along with their fast training, make them a compelling data-driven, black-box surrogate of dynamical systems, and a potential candidate for a component of digital twins.
@@ I Introduction
Echo state networks (ESN) and liquid state machines (LSM) were proposed as promising architectures for neural networks more than two decades ago. The ESN approach comes from the perspective of artificial neural networks using a discrete time formulation and sigmoid activation function [1]. The LSM approach, on the other hand, comes (more directly) from the perspective of biological neural networks, where the network vector field follows the integrate-and-fire neuron model in continuous time [2]. Soon after, the term reservoir was first invoked in Steil [3], and the term Reservoir Computing (RC) was coined in Verstraeten et al. [4] while attempting to implement these architectures in field-programmable gated arrays.
Before continuing further, we define some terms to avoid potential confusion. We use the term reservoir computing to mean the series of steps used to set up and train a recurrent neural network, which is treated as a driven dynamical system. The term reservoir refers to the driven state-space model; it has fixed or prescribed structure and has not yet been trained. The trained or autonomous reservoir refers to the autonomous state-space model, which incorporates the trained structure from a readout layer. As such, RC is a framework that utilizes the dynamics of an arbitrarily prescribed reservoir, where the training is performed only on the readout layer.
For the RC to have expressiveness in the space of nonlinear dynamics, either the reservoir dynamics or the readout must contain nonlinearity. Linear reservoir dynamics must be accompanied by a nonlinear readout, but nonlinear reservoir dynamics may be accompanied by either a linear or nonlinear readout. Most commonly, nonlinear reservoir dynamics are paired with a linear readout which is the case that this paper will focus on. RC with a nonlinear-nonlinear pairing has shown some success, but so far has not seen widespread usage [5].
The RC framework results in recurrent networks with some extremely promising aspects. It is able to predict autonomous dynamical systems with higher accuracy and smaller network sizes than deep feedforward networks and long-short term memory (LSTM) recurrent networks [6]. The training process is noniterative thus quick – if the dynamics are nonlinear, the training requires only a linear fit, compounding on the aforementioned benefits of RC. This is therefore a promising framework, with compelling reasons to understand its theoretical underpinnings and the source of its many strengths. This has yielded considerable research efforts since its conception, particularly in the role of contraction in the listening phase of RC [7, 8, 9, 10, 11].
@@
The recognition that GS and the ESP were closely related phenomena was a particularly fruitful realization [18, 19, 20, 21, 22] and one of the core motivations of this research effort. In this paper, we attempt to connect yet another concept – contraction theory [23, 24, 25, 26, 27] – with GS and the ESP. The ESP in continuous time reservoirs is relatively understudied compared to their discrete time counterparts. We give a preamble with discrete time reservoirs and ESP before pivoting to continuous time reservoirs. As far as the authors are aware of, Hart [20] provides the only results on the sufficient conditions of continuous time ESP. Our approach differs from that of Hart [20] but yields similar results that can be quickly computed.
The contraction that we invoke is better described as globally uniform state-contraction. Even though locally uniform state-contraction is sufficient for the ESP, the claims that we set out to make are nontrivial when considering all possible local cases, so we resort to the narrower global case. From here on, the prefix globally uniform state- is always implied when mentioning contraction, unless otherwise specified. It should also be understood that all inputs $u$ to the reservoir are measurements of an autonomous dynamical system with invertible velocity field and bounded trajectories, and that the description smooth will always mean continuously differentiable.
It is important to keep in mind that ESP and GS are only necessary conditions for the successful implementation of the RC framework; it guarantees very little about the resulting autonomous reservoir and the quality of its predictions. Even the universal approximation property (UAP) is not necessary, but a welcomed property – when possible – for a successfully trained reservoir to be a universal approximation of autonomous dynamical systems.
The organization of this paper is as follows. Section II goes over the basic structure of discrete time reservoirs. It gives the current understanding of the role of contraction and displays how certain contracting properties can be leveraged. Section III handles contraction in continuous time reservoirs, which bears significant resemblance to the approach in Section II. Our results provide a quickly computable sufficient condition that can be easily adjusted to guarantee contraction. Section IV discusses the UAP in the context of contracting activation functions, and topological conjugacy of the autonomous reservoir. Lastly, to avoid lengthy tangents, the appendices cover some specific details and claims made in the main sections.
@@ II Discrete Time Systems
$x_{t+1}=\sigma(Ax_{t}+Bu_{t}),$ (1)
$y_{t+1}=\sigma(Ay_{t}+Bu_{t}).$
These driven dynamical systems described in (1) are called reservoirs with $N$ nodes in the context of this paper. The connectivity matrix $A\in\mathbb{R}^{N\times N}$ describes the connections between the reservoir nodes. Matrix $B\in\mathbb{R}^{N\times M}$ is the input gain of the reservoir, designating how much each node is coupled to the inputs. The activation function, typically the hyperbolic tangent function, is the same for every node and is applied component-wise to its vector argument, i.e. $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{N}$ such that $\sigma([v_{1},v_{2}]^{\top})=[\sigma(v_{1}),\sigma(v_{2})]^{\top}$.
It is not expected that these two systems will reach a static equilibrium, since they are both driven by a time varying input $u$. Instead, consider defining a virtual displacement system (VDS) as $x_{t}-y_{t}$ to understand the long-term behavior of the individual systems [23]. Specifically, it is sufficient that if the VDS shrinks according to some norm for every $t\geq 0$, the two parent systems in (1) will asymptotically approach the same solution and be identically synchronized, which implies that the the reservoir is synchronized (in the sense of GS) to the dynamical systems generating the inputs [14, 16]:
$\norm{x_{t+1}-y_{t+1}}\leq k\norm{x_{t}-y_{t}}.$ (2)
For some vector $v$, let $\norm{v}$ denote any vector norm unless otherwise specified. The property described by (2) yields exponential stability with constant $10\leq k<$ 1, which also guarantees asymptotic stability. To demonstrate exponential stability, we make use of the contraction mapping property of the sigmoid family of functions for any two vectors $a\neq b$ in $\mathbb{R}^{N}$:
$\norm{\sigma(a)-\sigma(b)}\leq\norm{a-b}.$ (3)
In one dimension, contractions guarantee that two distinct points on its curve cannot generate a slope greater than unity and is equivalent to limiting the supremum magnitude of the derivative of $\sigma$ to unity. In higher dimensions, a contracting $\sigma$ can only shorten the distance between two vectors, even if a shift is applied. For some matrix $M$, let $\norm{M}$ denote any induced matrix norm. The matrix and vector norms throughout this paper will be implied through consistent usage of upper- and lower-case letters of the respective arguments. Using (3), we show that the VDS is globally exponentially stable when driven by any bounded $u$, so long as the induced matrix norm $\norm{A}<1$. See Appendix A for additional details.
$\norm{\sigma(Ax_{t}+Bu_{t})-\sigma(Ay_{t}+Bu_{t})}$ $\leq\norm{A(x_{t}-y_{t})}$ (4)
@@
Defining in this context $\norm{A}_{\ell}$ as some arbitrary norm, a broader sufficient condition for stability is $\min_{\ell}\norm{A}_{\ell}<1$. For ease of calculation, $\ell$ is usually chosen to correspond to the $1$-, $2$-, or $\infty$-norm. We have shown that when $\norm{A}<1$, the systems in (1) will uniformly and exponentially approach one another at a rate no slower than $\norm{A}$ regardless of initial conditions. If $A$ can be prescribed, then a reservoir can be constructed such that the VDS is always globally exponentially stable:
$\norm{x_{t}-y_{t}}\leq\norm{A}^{t}\norm{x_{0}-y_{0}}\Longrightarrow\lim_{t\to\infty}{x_{t}-y_{t}}=0.$ (5)
The rate at which the distance shrinks depend on the choice of norm. Also notice that the right-hand side of (4) is independent of $u_{t}$ and $B$. All this implies that a particular reservoir, defined by the triplet $(A,B,\sigma)$ with $\norm{A}<1$, will map every point of the input trajectory $u=\{u_{t}\}$ to a unique point in the reservoir trajectory $x=\{x_{t}\}$ in the infinite time limit [18, 14]. Necessarily, the ESP holds, and a unique GS occurs between the drive and response system:
$\lim_{t\to\infty}\frac{1}{t}\log\frac{\norm{x_{t}-y_{t}}}{\norm{x_{0}-y_{0}}}=\Lambda\leq\log\norm{A}<0.$ (6)
This condition on $A$ is closely related to the largest conditional Lyapunov exponent (CLE) of the system $\Lambda$, conditioned on the driving signal $u$. In fact, $\log\norm{A}$ is the negative-definite upper bound of $\Lambda$, which is itself the upper bound on all the other CLEs. This implies GS, so the reservoir implicitly has a globally attractive trajectory $x=\{x_{t}\}$ unique to any given $u=\{u_{t}\}$ as $t\to\infty$. In the infinite time limit, there exists a unique and smooth function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$, corresponding to the triplet $(A,B,\sigma)$ with constraint $\norm{A}\leq 1$, that maps the driving signal to the response, i.e. $x_{t}=\Phi(u_{t})$ as $t\to\infty$. The smoothness of $\Phi$ can be shown using Theorem 2.2.2 of Hart et al. [17]. For any desired error $\varepsilon>0$ and any chosen norm, there exist some function $\Phi$ and some time $T>0$ such that $\norm{x_{T}-\Phi(u_{T})}<\varepsilon$.
@@ III.1 Similarity to Discrete Time
In continuous time, the following situation is considered instead. Suppose that we are given some bounded and time varying input signal $u(t)\in\mathbb{R}^{M}$ in the time window $t\in[0,T]$ that is driving two identical reservoirs in the vector field according to the equations below:
$\dot{x}=-Cx+\sigma(Ax+Bu),$ (7)
$\dot{y}=-Cy+\sigma(Ay+Bu).$
As before, these two reservoirs have states $x(t)\in\mathbb{R}^{N}$ and $y(t)\in\mathbb{R}^{N}$ that stem from different initial conditions $x(0)\neq y(0)$. The matrices $A\in\mathbb{R}^{N\times N}$ and $B\in\mathbb{R}^{N\times M}$ serve the same purposes as the discrete time case, but an additional leak matrix $C\in\mathbb{R}^{N\times N}$ is necessary for stability, as will be demonstrated. The term leak is in reference to the leak current in biological neurons. For mathematical generality (as compared to biological realism), we impose that $C$ be symmetric positive definite (SPD). The activation function $\sigma$ is from the usual sigmoid family, applied component-wise as before. Define the VDS by $z(t)\equiv x(t)-y(t)$, with dynamics given by the following equation:
$\dot{z}=-C(x-y)+\sigma(Ax+Bu)-\sigma(Ay+Bu).$ (8)
@@
$\langle\dot{z},z\rangle=\underbrace{-\langle Cz,z\rangle}_{\leq-s_{N}(C)z^{2}}+\underbrace{\langle\sigma(Ax+Bu)-\sigma(Ay+Bu),x-y\rangle}_{\leq s_{1}(A)z^{2}}.$ (9)
@@
$\dot{V}(\dot{z},z)\leq-2kV(z)\Longrightarrow\norm{z(t)}_{2}\leq\norm{z(0)}_{2}e^{-kt}.$ (10)
Given the evolution of this Lyapunov function, $k>0$ results in the VDS being globally exponentially stable, making the parent systems (7) identically synchronized to one another, and the reservoir generally synchronized to the input. Since the evolution $\dot{V}(\dot{z},z)$ is ultimately a function of $z$ and $u$, $V(z)$, it is a control-Lyapunov function rather than a simple Lyapunov function. This is an important distinction as it suggests that $u$ plays a pivotal role in determining the attracting manifold, but plays no role determining whether the manifold is attracting. It is perhaps also a partial explanation as to why reservoirs robustly adapt to many input signals. Similar to the discrete time case, $-k$ is the upper bound of the largest CLE of the system:
$\lim_{t\to\infty}\frac{1}{t}\log\frac{\norm{z(t)}_{2}}{\norm{z(0)}_{2}}=\Lambda\leq-k<0.$ (11)
In the asymptotic limit, the attracting trajectory of $x(t)$ is unique for a given input $u(t)$, up to the quadruplet $(A,B,C,\sigma)$ and the aforementioned constraints on $A$ and $C$. It is common in practice to use the identity matrix $I$ instead of the generic SPD matrix $C$. Of course, the identity matrix is itself SPD, but this fixes $s_{N}(I)=1$ and restricts $s_{1}(A)=\norm{A}_{2}<1$, displaying contraction with same triplet $(A,B,\sigma)$ as the 2-norm discrete time case, but with leaky integrate-and-fire dynamics. In the infinite time limit, there exists a unique and smooth function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$ that maps $u(t)$ to $x(t)$, i.e. $x(t)=\Phi(u(t))$ as $t\to\infty$ [17, 28]. Though Hart et al. [17] handles the discrete time case, it makes use of contraction to show smoothness and uniqueness. We have shown that contraction occurs in the continuous time case, so $\Phi$ will similarly be smooth and unique. For any tolerance $\varepsilon>0$, there exist some function $\Phi$ and time $T>0$ such that $\norm{x(T)-\Phi(u(T))}_{2}<\varepsilon$.